OpenStudy (thunder33):

The graph of the equation below is a circle. What is the length of the radius of the circle? (x - 6)^2 + (y + 5)^2 = 15^2

OpenStudy (abmon98):

Standard form equation of a circle (x-h)^2+(y-k)^2=r^2 r^2=152 r=root(152)

OpenStudy (mjdennis):

So, if you are going to submit equations with squares, please use accurate notation.\[(x+a)^{2}\] should be typed in as (x+a)^2

OpenStudy (mjdennis):

I assume 152 is really \[15^{2}\] so you meant 15^2

OpenStudy (thunder33):

correct

OpenStudy (mjdennis):

Then did you get the correct answer by comparing @Abmon98 's form to your problem?